# Socialized Learning: A Survey of the Paradigm Shift for Edge Intelligence in Networked Systems

### v1

@@ VI-B2 Bidirectional Cooperation - Cooperation
The edge-cloud co-inference model, as shown in Fig. 29, exemplifies bidirectional cooperation in EI, with both clouds and edge devices actively participating, each contributing their unique strengths to enhance overall performance and efficiency.
  • Total Inference Latency Minimization: This collaborative dynamic optimizes task distribution based on the inherent capabilities of cloud and edge resources. The cloud, with its superior computational power, handles more complex, resource-intensive tasks, while edges, being closer to the data source, address latency-sensitive tasks. This judicious task distribution offloads certain operations to the cloud, allowing the edge to focus on timely processing, thereby enhancing system agility. Reinforcing this model, Li et al. [200] introduce AppealNet, an architecture aiming to synergize accuracy with low latency in inference. Similarly, Banitalebi-Dehkordi et al. [201] advance the collaborative edge-cloud AI through Auto-Split, a mechanism tailoring performance optimization with minimal latency.
  • Energy Consumption Minimization: Energy efficiency is another critical focus of the edge-cloud collaborative model. By intelligently offloading tasks to the cloud, edges can operate within their energy constraints, thereby promoting sustainable energy use across the EI system. Further, optimizing resource allocation and power management across this collaborative framework encapsulates the principle of mutual resource conservation. Xiang et al. [202] delve into the cooperative nuances of service provisioning in EC systems, advocating for an energy-efficient model. Concurrently, Zhang et al. [203] explore strategies to optimize energy consumption during task offloading.
@@
Harnessing both cloud and edge strengths ensures task processing with maximum efficiency, minimal latency, and energy conservation. However, due to challenges such as network latency, scalability, and data privacy, ongoing refinement and negotiation are essential to harness framework’s potential fully.
@@
  • Total Inference Latency Minimization: Analogous to social cooperation optimizing task delegation, the edge-device dynamic leverages the strengths of both tiers for latency reduction. Edges, equipped with robust computational resources, absorb significant processing loads, enabling devices to efficiently manage local tasks and circumvent excessive transmission delays. Such an interaction finds support in the work of Bai et al. [204], where the balance between benefits of computational offloading and challenges of communication latency is assessed, highlighting the utility of intelligent reflecting surfaces (IRSs) in enhancing efficiency.
  • Energy Consumption Minimization: Drawing parallels with social emphasis on resource preservation, the edge-device collaboration prioritizes energy conservation. Distributed computational duties and focused energy strategies, as seen in Krouka et al. [205], implement model compression techniques for energy optimization at the edge without accuracy compromise. Likewise, Xiao et al. [206] propose a dynamic, RL-driven approach, aligning devices to real-time conditions for sustained energy efficiency.
@@
where $L$ and $E$ are the total inference latency and energy consumption, respectively. $L_{e}(T_{e})$ and $L_{d}(T_{d})$ are the latency of executing tasks $T_{e}$ and $T_{d}$ on the edge and device, respectively. $E_{e}(T_{e})$ and $E_{d}(T_{d})$ are the energy consumption of executing tasks $T_{e}$ and $T_{d}$ on the edge and device, respectively. $T_{e}$ and $T_{d}$ are the proportions of tasks executed on the edge and device, respectively. Building on edge-cloud co-inference, edge-device co-inference highlights bidirectional cooperation. Emphasizing task offloading, adaptive management, and proximity benefits, this mechanism minimizes latency and ensures energy efficiency.
Figure 31: Illustration of Decision Interaction.
@@
  • Total Inference Latency Minimization: Bidirectional decision-making cooperation aims to minimize inference latency through efficient knowledge pooling. This is achieved by fostering a cooperative environment where devices exchange decision-related insights or intermediate results, synergizing their expertise for timely decisions. An example is presented by Carroll et al. [207], where the Flexi BiT framework illuminates the potential of masked modeling in sequential decision-making. Just as collective decision-making in societies often taps into diverse knowledge sources for better outcomes, this framework’s capacity to be attuned to a range of tasks exemplifies the benefits of collaborative decision-making in the computational realm.
  • Energy Consumption Minimization: In the bidirectional decision-making domain, devices are keenly aware of energy saving. They engage in strategies like intelligent resource allocation or the selective sharing of decision data, all aimed at optimizing energy consumption without compromising decision accuracy. Tang et al. [208] present an exploration into the optimization of energy in distributed inference, consistent with the social principle of strategic resource division for the collective good.
@@ VI-B3 Bidirectional Cooperation - Exchange
In social interactions, exchanging resources and knowledge is crucial for mutual benefit. This principle, when applied to EI, entails an interaction between hierarchical layers during the inference phase. The cloud, with its extensive data repositories, requires data from edges for context-aware inferences, while edge layers use the cloud’s computational power for deeper analysis. This interplay ensures inferences are prompt and relevant, and capitalize on each stratum’s strengths, fostering an integrated and efficient system. We explore how clouds, edges, and devices while maintaining their distinct objectives, engage in this reciprocal exchange during task inference.
@@
  • Energy Consumption Minimization: The principle of bidirectional exchange also applies to energy consumption. Each layer aims to use energy efficiently in making inferences. The cloud leverages its vast resources, offloading tasks to the edge to conserve energy. The edge, due to its proximity to data sources, minimizes energy expenditure during data transmission and focuses on localized inferences. Devices, inherently resource-constrained, advocate for energy-efficient inferences and selective data transmissions. Hu et al. [210] spotlight energy conservation in mobile networks, illustrating how edge situated closer to mobile devices can optimize energy consumption. When a device’s computational resources are insufficient, offloading tasks to the edge presents an energy-conservative solution.
[…]
$\underset{X_{C},X_{E},X_{D}}{\text{max}}$ $\quad\theta_{C}O_{C}(X_{C})+\theta_{E}O_{E}(X_{E})+\theta_{D}O_{D}(X_{D})$ (30)
$-\lambda[L_{c}(X_{C})+L_{e}(X_{E})+L_{d}(X_{D})]$
$-\omega[E_{c}(X_{C})+E_{e}(X_{E})+E_{d}(X_{D})],$
s.t. $\quad X_{C}+X_{E}+X_{D}\leq T.$
Task and data exchanges between the cloud ($X_{C}$), edge ($X_{E}$), and devices ($X_{D}$) are optimized in our model to meet the objectives of each ($O_{C}$, $O_{E}$, $O_{D}$), while considering latency ($L_{c}$, $L_{e}$, $L_{d}$) and energy consumption ($E_{c}$, $E_{e}$, $E_{d}$). Weights ($\theta_{C}$, $\theta_{E}$, $\theta_{D}$, $\lambda$, $\omega$) prioritize these factors accordingly, ensuring that the total exchange ($T$) aligns with system capabilities and objectives. This inter-layer collaboration, mirroring social reciprocity, improves task allocation, reduces latency and conserves energy by leveraging participant’s unique strengths.
The discussion on Socialized Inference in EI draws an analogy between social cooperation and collaborative decision-making, such as edge-device and cloud-edge interactions. Detailed in Table XI, these approaches manage the trade-off between latency and energy efficiency, some focusing on real-time response and others on energy savings.
VI-C Intra-layer Submissive Interaction in Socialized Inference
VI-C1 Conformity
In EI, as in society, devices align with norms for cohesion and efficiency.
@@ VI-C1 Conformity
Conformity in EI focuses on enabling ultra-reliable and low-latency communication (URLLC) by streamlining processes and eliminating superfluous exchanges to facilitate swift and reliable decision-making. To encapsulate these ideas into a formula, we can represent the system’s efficiency as a function of conformity, latency, and reliability. The SL function Eq. (1) for ultra-reliable and low-latency communication is $F([\text{devices}],[\text{conformity}],\varnothing,\text{inference},\text{tasks})$. The optimization objective is:
@@
$\underset{f,C}{\text{max}}\quad C\times f-\alpha\times L,\quad\text{s.t.}\quad 0\leq C,f\leq 1,L\geq 0,$ (32)
@@ VI-D1 Unidirectional Submissive Interaction - Suggestion&imitation
In traditional societies, authoritative figures guide behaviors. The suggestion-imitation pattern, depicted in Fig. 33, shows how leaders guide actions. In EI inference, layers with contextual or computational advantages provide ‘suggestions’ as processed insights. Other layers then ‘imitate’ by aligning their strategies with these insights. Unlike model parameter and feature transfer in Socialized Training, Socialized Inference focuses on real-time decision alignment, harmonizing actions with guiding layers. This ensures EI cohesion, reflecting how authoritative figures influence group decisions. We show how different layers’ dynamics affect communication latency, energy usage, and computation.
  • Resource Allocation: Resource allocation in EI, emulating social dynamics, involves the harnessing of resources and the imitation of efficient strategies. Authoritative layers, like a cloud or primary controller, guide resource allocation by suggesting optimal routes or efficient edge strategies. Their authority enhances effectiveness, steering lower layers towards optimization, thus reducing communication delays. on data routing or task allocation minimize back-and-forth communication, reducing overhead. Cen et al. [214] explore a reinforcement learning approach to these challenges in cloud-edge environments. Similarly, Hu et al. [215] investigate intelligent resource allocation in edge-cloud networks using a hybrid DDPG-D3QN model.
  • Energy Management: Authoritative layers, through their suggestions, can drive energy-efficient strategies, urging edges to mimic energy-saving techniques or adopt operational modes. Such imitation dynamics prolong device lifespans, reducing overall energy wastage. Ruan et al. [216] proffer a three-tier architecture, emphasizing latency improvements and efficient energy scheduling.
  • Computational Offloading: Task balancing between cloud and edge is central to EI. Authoritative entities steer edge devices’ offloading decisions for balanced computational loads. This approach also reduces communication latency by offloading to closer edges and adopting efficient practices, enhancing system responsiveness. Wang et al. [217] focus on computational offloading in wireless networks, balancing edge server capacity and energy use. Zeng et al.[218] present a joint optimization for computational offloading, assessing its effects on energy use and performance.
@@
$\underset{S,I}{\text{max}}$ $\quad\theta_{1}R(S,I)+\theta_{2}E(S,I)+\theta_{3}O(S,I),$ (33)
@@
Unidirectional submissive interactions, like suggestion and imitation dynamics, are evident in both human societies and EI inference. The parallels drawn between social interactions and the inference processes in EI emphasize the potential benefits of such collaborative frameworks. Embracing these dynamics enables a harmonized, efficient, and responsive system, optimizing resource allocation, energy management, and computational tasks. This hierarchical structure optimizes processes from resource allocation to energy management. A comparative analysis presented in Table XII elucidates interactions’ distinctions and synergies.
@@ VI-E1 Competition
  • Resource Allocation: Intra-layer competition centers on resource allocation. Devices competing for limited computational resources like processing power or storage may increase communication demands through resource negotiation and adversely affect communication latency. Additionally, increased competition often leads to higher energy demands, compromising system-wide energy efficiency. Fan et al. [165] investigate networks where agents compete for resources, proposing a mechanism for efficient and fair allocation. Sun et al. [220] optimize for maximum network satisfaction through fair resource sharing.
  • Task Offloading: In edge-cloud collaborations, task offloading becomes another aspect of intra-layer competition. Devices jostle for offloading their tasks to available, capable nodes, potentially affecting latency and energy consumption. Offloading strategies can either promote energy efficiency or, if unchecked, exacerbate energy demands. Ji et al. [221] contribute to this discourse by focusing on energy-efficient offloading algorithms, considering latency tolerances and uncertainties inherent in such networks.
  • Computing Efficiency: The competition also invigorates devices to bolster their computing efficiency. Devices vie to hone algorithms, employ cutting-EC methods, or adopt strategies that promise efficiency. Such pursuits, while fostering innovation, could inadvertently augment computational overheads. Lim et al. [222] provide insight into scheduling algorithms in allocation scenarios and offer effective strategies for task execution in competitive environments. Zhang et al. [223] expand on this by talking about computing schedules in EI that are energy-efficient.
@@ VI-F1 Unidirectional Competition - Compulsory
In social hierarchies, upper echelons dictate the actions of subordinate layers, establishing a unidirectional command chain, as depicted in Fig. 35. Although this hierarchy centralizes decision-making and optimizes inference, it also engenders issues such as communication delays, energy demands, and computing inefficiency. Mitigating these challenges necessitates collaborative strategies in resource allocation, task offloading, and scheduling, thereby enhancing resource efficiency, communication efficacy, and overall performance.
@@
  • Task Offloading: In the edge-cloud collaboration, unidirectional compulsion is manifest when superior layers necessitate task offloading to their lower counterparts. This mechanism enhances centralized management and scalability but also raises latency issues. Almutairi et al. [63] delve into task offloading, highlighting a fuzzy logic system to optimize the end-to-end service time. Proximity-based communication and energy-efficient offloading ensure tasks find their way to energy-optimized nodes.
  • Computing Efficiency: Unidirectional competition within layers during inference accentuates the edge-cloud collaborative system’s computing efficiency. Although this competition can be demanding, it also catalyzes advancements in performance. Given the heterogeneity of computational resources, collaborative algorithms and ML approaches are needed. Maray et al. [226] introduce a dynamic time-sensitive scheduling algorithm, while Fan et al. [227] aim at optimizing the total task processing delay. Furthermore, Zhang et al. [228] bring forth a mechanism optimizing the task completion time, emphasizing the limited storage space of the ES and its implications.
@@
[…]

### v2

@@ VI-B2 Bidirectional Cooperation - Cooperation
The edge-cloud co-inference, as shown in Fig. 29, exemplifies bidirectional cooperation in EI. In this scheme, both cloud and edge devices actively participate, leveraging their unique strengths to enhance overall performance and efficiency.
  • Total Inference Latency Minimization: This collaborative dynamic optimizes task allocation based on the inherent capabilities of cloud and edge resources. The cloud, with its superior computing power, handles more complex, resource-intensive tasks, while edges, being closer to the data source, address latency-sensitive tasks. This judicious task allocation offloads certain operations to the cloud, allowing the edge to focus on timely processing, thereby enhancing system adaption. Reinforcing this scheme, Li et al. [203] introduce AppealNet, an architecture aiming to synergize accuracy with low latency in inference. Similarly, Banitalebi Dehkordi et al. [204] advance the collaborative edge-cloud AI through Auto-Split, a mechanism tailoring performance optimization with minimum latency.
  • Energy Consumption Minimization: Energy efficiency is another critical focus of the edge-cloud co-inference. By intelligently offloading tasks to the cloud, edges can operate within their energy constraints, thereby promoting sustainable energy use across the EI system. Further, optimizing resource allocation and power management across this collaborative framework encapsulates the principle of mutual resource conservation. Xiang et al. [205] delve into the cooperation of service provisioning in EC systems, advocating for an energy-efficient scheme. Concurrently, Zhang et al. [206] explore strategies to optimize energy consumption during task offloading in edge-cloud co-inference.
@@
Harnessing both cloud and edge strengths ensures task processing with maximum efficiency, minimum latency and energy conservation. Due to challenges such as network latency, scalability and data privacy, ongoing refinement and adjustment is essential to harness the framework’s potential.
@@
  • Total Inference Latency Minimization: Analogous to social cooperation optimizing task offloading, the edge-device dynamic leverages the strengths of both tiers for latency reduction. Edges, equipped with robust computational resources, absorb significant processing loads, enabling devices to efficiently manage local tasks and circumvent excessive transmission delays. Such an interaction finds support in the work of Bai et al. [207], where the balance between benefits of computational offloading and challenges of communication latency is assessed, highlighting the utility of intelligent reflecting surfaces (IRSs) in enhancing efficiency.
  • Energy Consumption Minimization: Drawing parallels with social emphasis on resource preservation, the edge-device collaboration prioritizes energy conservation. Distributed energy strategies, as seen in Krouka et al. [208], implement model compression techniques for energy optimization at the edge without model performance compromise. Likewise, Xiao et al. [209] propose a dynamic, RL-driven approach, aligning devices to real-time conditions for sustained energy efficiency.
@@
where $L$ and $E$ are the total inference latency and energy consumption, respectively. $L_{e}(T_{e})$ and $L_{d}(T_{d})$ are the latency of executing tasks $T_{e}$ and $T_{d}$ on the edge and device, respectively. $E_{e}(T_{e})$ and $E_{d}(T_{d})$ are the energy consumption of executing tasks $T_{e}$ and $T_{d}$ on the edge and device, respectively. $T_{e}$ and $T_{d}$ are the proportions of tasks executed on the edge and device, respectively. Building on edge-cloud co-inference, edge-device co-inference greater highlights the benefits of bidirectional cooperation. Emphasizing task offloading, adaptive management, and proximity benefits, this mechanism minimizes latency and ensures energy efficiency.
Figure 31: Illustration of decision interaction in socialized inference.
@@
  • Total Inference Latency Minimization: Bidirectional decision-making cooperation aims to minimize inference latency through efficient knowledge pooling. This is achieved by fostering a cooperative environment where devices exchange decision-related insights or intermediate results, synergizing their expertise for timely decisions. An example is presented by Carroll et al. [210], where the FlexiBiT framework illuminates the potential of masked modeling in sequential decision-making. Just as collective decision-making in societies often utilises diverse knowledge sources for better outcomes, this framework’s capacity to be attuned to a range of tasks exemplifies the benefits of collaborative decision-making.
  • Energy Consumption Minimization: In the bidirectional decision-making domain, devices are keenly aware of energy saving. They engage in strategies like intelligent resource allocation or the selective sharing of decision data, all aimed at optimizing energy consumption without compromising decision accuracy. Guo et al. [211] present an exploration into the optimization of energy consumption in distributed inference, consistent with the social principle of strategic resource division for the collective good.
@@ VI-B3 Bidirectional Cooperation - Exchange
In social interactions, exchanging resources and knowledge is crucial for mutual benefit. This principle, when applied to EI, entails an interaction between hierarchical layers during the inference phase. The cloud, with its extensive data repositories, requires data from edges for context-aware inferences, while edge layers use the cloud’s computing power for deeper analysis. This interplay ensures inferences are prompt and relevant, and capitalize on each stratum’s strengths, fostering an integrated and efficient system. We explore how clouds, edges, and devices while maintaining their distinct objectives, engage in this reciprocal exchange during task inference.
@@
  • Energy Consumption Minimization: The principle of bidirectional exchange also applies to energy consumption. Each layer aims to use energy efficiently in making inferences. The cloud leverages its vast resources, offloading tasks to the edge to conserve energy. The edge, due to its proximity to data sources, minimizes energy consumption during data transmission and focuses on local inferences. Devices, inherently resource-constrained, advocate for energy-efficient inferences and selective data transmissions. Hu et al. [213] spotlight energy conservation in mobile networks, illustrating how edge situated closer to mobile devices can optimize energy consumption. When a device’s computational resources are insufficient, offloading tasks to the edge presents an energy-conservative solution.
[…]
max $\quad\theta_{C}O_{C}(X_{C})+\theta_{E}O_{E}(X_{E})+\theta_{D}O_{D}(X_{D})$ (30)
$-\lambda[L_{c}(X_{C})+L_{e}(X_{E})+L_{d}(X_{D})]$
$-\omega[E_{c}(X_{C})+E_{e}(X_{E})+E_{d}(X_{D})],$
s.t. $\quad X_{C}+X_{E}+X_{D}\leq T.$
Task and data exchanges between the cloud ($X_{C}$), edge ($X_{E}$), and devices ($X_{D}$) are optimized in our model to meet the objectives of each ($O_{C}$, $O_{E}$, $O_{D}$), while considering latency ($L_{c}$, $L_{e}$, $L_{d}$) and energy consumption ($E_{c}$, $E_{e}$, $E_{d}$). Weights ($\theta_{C}$, $\theta_{E}$, $\theta_{D}$, $\lambda$, $\omega$) prioritize these factors accordingly, ensuring that the total exchange ($T$) aligns with system capabilities and objectives. This inter-layer collaboration, mirroring social reciprocity, improves task allocation efficiency, reduces latency and conserves energy by leveraging participant’s strengths.
The discussion on Socialized Inference in EI draws an analogy between social cooperation and collaborative decision-making, such as edge-device and cloud-edge interactions. Detailed in Table XI, these mechanisms manage the trade-off between latency and energy efficiency, some focusing on real-time response and others on energy savings.
VI-C Intra-layer Submissive Interaction in Socialized Inference
VI-C1 Conformity
Similar to how individuals in society align their actions with group norms to promote cohesion, devices also adhere to established norms to enhance efficiency.
@@ VI-C1 Conformity
Conformity in EI focuses on enabling ultra-reliable and low-latency communication (URLLC) by streamlining processes and eliminating superfluous exchanges to facilitate swift and reliable decision-making. To encapsulate these ideas into a formula, we can represent the system’s efficiency as a function of conformity, latency, and reliability. The SL function Eq. (1) for URLLC is $F([\text{devices}],[\text{conformity}],\varnothing,\text{inference},\text{tasks})$. The optimization objective is:
@@
${\text{max}}\quad C\times f-\alpha\times L,\quad\text{s.t.}\quad 0\leq C,f\leq 1,L\geq 0,$ (32)
@@ VI-D1 Unidirectional Submissive Interaction - Suggestion&imitation
In traditional societies, authoritative figures guide behaviors. The suggestion&imitation pattern, depicted in Fig. 33, shows how leaders guide actions. In EI inference, layers with contextual or computational advantages provide ‘suggestions’ as processed insights. Other layers then ‘imitate’ by aligning their strategies with these insights. Unlike model parameter and feature transfer in Socialized Training, Socialized Inference focuses on real-time decision alignment, harmonizing actions with guiding layers. This ensures EI cohesion, reflecting how authoritative figures influence group decisions. We show how different layers’ dynamics affect communication latency, energy usage, and computation.
  • Resource Allocation: Resource allocation in EI, emulating social dynamics, involves the harnessing of resources and the imitation of efficient strategies. Authoritative layers guide resource allocation by suggesting optimal routes or efficient edge strategies, steering lower layers towards optimization and reducing communication delays. Cen et al. [217] explore an RL approach to these challenges in cloud-edge environments. Similarly, Hu et al. [218] investigate intelligent resource allocation in edge-cloud networks using a hybrid DDPG-D3QN model.
  • Energy Management: Authoritative layers’ suggestions drive energy-efficient strategies, urging edges to mimic energy-saving techniques or adopt operational modes. Such imitation dynamics prolong device lifespans, reducing overall energy wastage. Ruan et al. [219] propose a three-tier architecture, emphasizing latency improvements and energy-efficient scheduling.
  • Computational Offloading: Task balancing between cloud and edge is central to EI. Authoritative entities steer lower layers’ offloading decisions for balanced computational loads, reducing communication latency and enhancing system responsiveness. Wang et al. [220] focus on computational offloading in wireless networks, balancing ES capacity and energy consumption. Zeng et al. [221] present a joint optimization for computational offloading, assessing its effects on energy usage and performance.
@@
max $\quad\theta_{1}R(S,I)+\theta_{2}E(S,I)+\theta_{3}O(S,I),$ (33)
@@
Unidirectional submissive interactions, like suggestion and imitation dynamics, are evident in both human societies and EI inference, emphasizing the potential benefits of such collaborative frameworks. Embracing these dynamics enables a harmonized, efficient, and responsive system, optimizing resource allocation, energy management, and computational tasks. This hierarchical structure optimizes processes from resource allocation to energy management. A comparative analysis presented in Table XII elucidates interactions’ distinctions and synergies.
@@ VI-E1 Competition
  • Resource Allocation: Intra-layer competition centers on resource allocation. Devices competing for limited computational resources may increase communication demands, adversely affecting communication latency. Additionally, increased competition often leads to higher energy demands, compromising the energy efficiency of the system. Fan et al. [168] investigate networks where agents compete for resources, proposing a mechanism for efficient and fair allocation. Sun et al. [223] optimize for maximum network satisfaction through fair resource sharing.
  • Task Offloading: In edge-cloud collaborations, task offloading becomes another aspect of intra-layer competition. Devices jostle for offloading their tasks to available, capable nodes, potentially affecting latency and energy consumption. Offloading strategies can either promote energy efficiency or, if unchecked, exacerbate energy demands. Ji et al. [224] contribute to this issue by focusing on energy-efficient offloading algorithms, considering latency tolerances and uncertainties inherent in such networks.
  • Computing Efficiency: The competition is also driving devices to bolster their computing efficiency. Devices vie to optimize algorithms, employ cutting-EC methods, or adopt strategies that promise efficiency. Such pursuits, while fostering innovation, could inadvertently increase computational overheads. Lim et al. [225] provide insight into scheduling algorithms in allocation scenarios and offer effective strategies for task execution in competitive environments. Zhang et al. [226] expand on this by discussing computing schedules in EI that are energy-efficient.
@@ VI-F1 Unidirectional Competition - Compulsory
In social hierarchies, upper echelons dictate the actions of subordinate layers, establishing a unidirectional command chain, as depicted in Fig. 35. Although this hierarchy centralizes decision-making and optimizes inference, it also generates issues such as communication delays, energy demands, and computing inefficiency. Mitigating these challenges necessitates collaborative strategies in resource allocation, task offloading, and scheduling, thereby enhancing resource efficiency, communication efficiency, and overall performance.
@@
  • Task Offloading: In the edge-cloud collaboration, unidirectional compulsory is manifest when superior layers necessitate task offloading to their lower counterparts. This mechanism enhances centralized management and scalability but also raises latency issues. Almutairi et al. [66] delve into task offloading, highlighting a fuzzy logic system to optimize the end-to-end service time. Proximity-based communication and energy-efficient offloading ensure tasks find their way to energy-optimized nodes.
  • Computing Efficiency: Unidirectional compulsory within layers during inference accentuates the edge-cloud collaborative system’s computing efficiency. Although this compulsory can be demanding, it also catalyzes advancements in performance. Given the heterogeneity of computational resources, collaborative algorithms and ML approaches are needed. Maray et al. [229] introduce a dynamic time-sensitive scheduling algorithm, while Fan et al. [230] aim at optimizing the total task processing delay. Furthermore, Zhang et al. [231] bring forth a mechanism optimizing the task completion time, emphasizing the limited storage space of the ES and its implications.
@@
max $\quad\delta_{1}\cdot R(H)+\delta_{2}\cdot T(H)+\delta_{3}\cdot Comp(H),$ (35)